OpenStudy (anonymous):

What is the solution to 1,9647 • 102?

OpenStudy (anonymous):

http://www.online-calculator.com/

OpenStudy (anonymous):

i already have a calculator but i cant do a 10^2 on the calculator but thx for trying to help :)

OpenStudy (anonymous):

You multiply 10*10

OpenStudy (anonymous):

Then multiply is by 1.96

OpenStudy (anonymous):

10^2 = 10 * 10

OpenStudy (anonymous):

*it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

10^3= 10 * 10 * 10 10^4= 10 * 10 * 10 * 10